# Inflation with vector fields revisited: heavy entropy perturbations and primordial black holes

### v2

@@ 2.3 Evolution of the adiabatic mode
where $N_{c}$ is the time of horizon-crossing, $N_{e}$ is the time of end of inflation and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing. For small $h$ this is a small correction to the adiabatic mode. The zero-th order entropy perturbation is decoupled with the adiabatic one and massless in the slow-roll limit. Hence the zeroth-order power spectrum of entropy perturbation is simple $\mathcal{P}_{\mathcal{F}}^{(0)}=H^{2}/(4\pi^{2})$. Then the leading-order correction to the power spectrum of curvature perturbation can be estimated to
@@ 3 EFT under massive entropy mode
3.1 Modified dispersion relation
From the quadratic Lagrangian (2.2) we can write down the linear equations of motion
@@ 3.1 Modified dispersion relation
$\ddot{\zeta_{c}}+p^{2}\zeta_{c}=$ $\sqrt{2}hH\mathcal{A}\dot{\mathcal{F}},$
¨$\ddot{\mathcal{F}}+p^{2}\mathcal{F}+m_{s}^{2}\mathcal{F}=$ $-\sqrt{2}hH\mathcal{A}\dot{\zeta}_{c},$ (3.2)
$\zeta_{c}=\zeta_{+}e^{i\omega_{+}t}+\zeta_{-}e^{i\omega_{-}t},$
$\mathcal{F}=\mathcal{F}_{+}e^{i\omega_{+}t}+\mathcal{F}_{-}e^{i\omega_{-}t},$ (3.3)
where the two frequencies $\omega_{\pm}$ are given by
$\omega_{\pm}^{2}=\frac{m_{s}^{2}}{2c_{s}^{2}}+p^{2}\pm\frac{m_{s}^{2}}{2c_{s}^{2}}\sqrt{1+\frac{4\left(1-c_{s}^{2}\right)p^{2}}{m_{s}^{2}c_{s}^{-2}}}$ (3.4)
and the speed of sound is given by
$\frac{1}{c_{s}^{2}}=1+\frac{32H^{2}h^{2}}{m_{s}^{2}}.$ (3.5)
The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$. The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$, which requires $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$. Hence we have a cut-off scale of EFT (B.8) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale. The UV scale is given by
$p_{\text{UV}}^{2}=\frac{|m_{s}^{2}c_{s}^{-2}|}{4\left(1-c_{s}^{2}\right)}.$ (3.6)
3.2 Integrating out the heavy $\mathcal{F}$
The fluctuation is heavy for $h\gg 1$ and we also have $|m_{s}^{2}|\gg m_{\sigma}^{2}$ as we have mentioned before. For positive $m_{s}^{2}$, the operator from coupling $\dot{\zeta}\mathcal{F}$ would reduce the speed of sound of the fluctuation [43]. However from (2.20) the mass square is negative due to the contribution from kinetic coupling $f(\phi)$. We can ignored the slow-roll suppressed terms and the entropy mass is given by (2.20). Then the speed of sound in this decoupling limit is
$\frac{1}{c_{s}^{2}}=1-\frac{4\left(1+2h^{2}\right)}{3+2h^{2}}.$ (3.7)
The speed of sound $c_{s}^{2}\simeq-1/3$ is imaginary for heavy entropy mass $h\gg 1$. The imaginary speed of sound has been studied in the two-scalar inflation recently [12, 14, 44]. And the entropy mass-square in the two-scalar models is also imaginary hence is tachyonic before the sound horizon crossing $p^{2}c_{s}^{2}\sim H^{2}$. The curvature fluctuation sourced by the entropy fluctuation also experiences transient instability before horizon crossing. After the horizon-crossing, the dynamics of fluctuations are influenced by the background. In other words, the mass term $m_{\sigma}$ and the Hubble terms become important outside the horizon.
Figure 1: Evolution of entropy perturbation $\mathcal{F}$ for $h=10$. An instability before horizon-crossing and constant evolution on super-horizon are presented. The scales are arbitrary.
We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime (see Appendix B for reviews). In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator
@@ 3.2 Integrating out the heavy $\mathcal{F}$
Then the leading-order of the linear equation of motion of $\mathcal{F}$ is given by
@@
We show the consistent of the leading-order contribution of $\mathcal{F}$ in Figure 1 for large enough $h$. The evolution of curvature perturbation on large scales is sourced by non-adiabatic perturbation $\dot{\zeta}=-\delta P_{\text{nad}}H/(\rho+p)$ [45, 46, 47]. Then from the leading-order solution of entropy perturbation, the non-adiabatic mode can be represented as
The adiabatic initial condition ensures that the r.h.s. should be relatively small enough on large scales. We can see that the entropy fluctuation doesn’t decay outside the horizon, but is proportional to the curvature fluctuation if $\zeta=\text{constant}$ on super-horizon.
We can find that in the EFT of single field, we have only this constant curvature perturbation on large scales. In the slow-roll limit the entropy mass $m_{s}$ can be calculated by the (2.20). From (3.6) the EFT is only valid under scale
@@
where $c_{s}$ is given by (3.7). We have used the integration by parts to eliminate the $\dot{\zeta}\zeta$ terms in the action. Then we found that the mass term of curvature fluctuation is cancelled and $\zeta$ is conserved on super-horizon scales.
Figure 2: The adiabaticity of super-horizon ($|k\tau_{e}|=10^{-15}$) curvature perturbation. As $h$ increases from very small values, the perturbations become non-adiabatic. However, when $h$ exceeds $\mathcal{O}(0.1)$, the perturbation quickly becomes adiabatic.
Different from multi-scalar inflation, where the non-decaying non-adibatic modes will source to curvature modes so that $\zeta$ evolves on super-horizon. In dilation-gauge inflation, we have constant $\zeta$ for non-decaying $\mathcal{F}$. In other words, $\mathcal{F}$ doesn’t play the role of non-adiabatic mode in dilaton-gauge inflation. For mid-value $h\lesssim\mathcal{O}(0.1)$, where $|m_{s}|\sim H$, the sub-leading terms in $\delta P_{\text{nad}}$ become important and the curvature perturbation evolves on super-horizon. These non-adiabatic contributions are from the cumulative modes (2.27). We show this non-adiabatic occurrence for mild $h\lesssim\mathcal{O}(0.1)$ in Figure 2.
One of the most important findings from this result is that the cumulative modes (2.27) decay away for large $h$. Because for large $h$, the modes (3.9) on super-horizon are exactly the constant solutions of the adiabatic mode (2.23). The decaying modes can be also obtained through solving the equation of motion (3.1), which has been studied in spatially flat gauge in [20]. The decaying mode is $\zeta\propto(-k\tau)^{(3-\sqrt{9-96h^{2}})/2}$. For small $h$ we can expand this solution and find that the leading correction is proportional to e-folding number $\ln(-k\tau)=N$, which is the solution (2.27).
@@ 4 Primordial black holes
The remained adiabatic mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound. Therefore the power spectrum is enhanced exponentially. If vector fields switch on only at late time, The power spectrum is enhanced on small scales so the formations of PBHs in radiation era is possible [48, 49]. And these PBHs can be one of candidates of the dark matter [50, 51, 52] and supermassive black holes in our universe [53]. The enhancement of the power spectrum due to the heavy fields is studied recently in multiple scalar-field models, and is used to discuss the PBHs formation [56, 54, 55]. In previous section, we have realized that it also can be achieved by using vector fields.
We approach to the PBHs without any specific models ($V(\phi)$ and $f(\phi)$). We here only consider the switch-on of the vector fields at late time $N_{f}$ during inflation and see the features of power spectrum on small scales for generating PBHs. We choose the horizon-crossing time of the CMB scale $k_{\text{CMB}}=0.05\text{Mpc}^{-1}$ as $N_{\text{CMB}}=0$. Assuming that $h$ has peak at $N_{f}$. The curvature modes with different scales that cross the horizon at different time have different contributions to the power spectrum. We can define the corresponding scale
@@
The modes $k\gg k_{f}$ that cross the horizon far after the switch-off time experienced nothing hence the power spectrum is the same as the single-field inflation $\mathcal{P}_{\zeta}^{(0)}=H^{2}/(8\pi\epsilon M_{\text{pl}}^{2})$ on small scales.
On the other hand, the modes $k\ll k_{f}$ that cross the horizon far before the switch-on time will be impacted by the entropy fluctuation on super-horizon. The entropy fluctuation doesn’t decay outside the horizon hence always transfer to the curvature fluctuation. Hence on these large scales, the power spectrum is shifted so different from the single-field one.
The interesting scales are the modes $k\sim k_{f}$ that cross the horizon near the switch-on time $N_{f}$ have large enhancements due to the large instability of the entropy fluctuation. Hence we expect that the power spectrum of the curvature fluctuation has exponential enhancements near the scale $k_{f}$. We can see from Figure 1 that the mode takes a few e-folding to grow before frozen-in outside the horizon. Hence there are two kinds of power spectrum, the broad ($\delta N\gtrsim\ln{(h_{0})}$) and the sharp ($\delta N\lesssim\ln{(h_{0})}$) switch-on of $h(N)$ [56]. For the shape case, the curvature modes stop growing before crossing the horizon.
4.1 Board case
For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{v}$ satisfying
@@ 4.1 Board case
Then the most contributions to the power spectrum are the modes with $N_{v}\sim N_{f}$. We here consider the Gaussian profile of the function
@@
where the profile is controlled by two parameters $(h_{0},\delta N)$. On super-horizon, the entropy modes do not decay and still couple to the curvature ones. Hence we should discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
@@
To characterize the peak of power spectrum with these modes, it’s also useful to introduce a dimensionless parameter $\mathcal{I}$ [44, 57]. The quantization of the curvature fluctuation with imaginary speed of sound is quite different from the real one [44]. The difference is that unlike the Bunch-Davies vacuum, the negative frequency modes are non-zero for initial state of EFT due to the different commutation conditions of the mode functions. Taking the super-horizon limit, the leading contribution of the power spectrum is
@@
For modes who start to experiences transient instability earlier, the amplitude will be larger after crossing the horizon. And we need to use a more UV EFT to describe these modes on deeper sub-horizon scales.
Figure 3: The numerical and analytical results (4.4) of power spectrum for the Gaussian case, where $h_{0}=8$ and $\delta N=2$.
The scale of peak $k_{p}$ is the one that has maximal $h$ at time when the modes is growing at $N_{v}$, i.e., $N_{p_{k}}=N_{v}$. Then we have
@@
We show the numerical and analytical results of our calculation in the Figure 3. One thing we should mention is that (4.4) is not valid for all range of $h(N)$ because the computation is under the vaild of EFT so the mass term of curvature is small compared to the heavy entropy one (or other, large $h(N)$). Hence this estimation work more well for $|m_{s}^{2}(N)|\gtrsim 8H^{2}$ in our Gaussian-profile model.
@@
As Figure 3 shown, for the l.h.s of the peak, where the switch occurs after crossing the horizon, the power spectrum is governed by the long range contributions. Hence we can see a shift on the large scale due to the coupling with entropy fluctuation on large scales. We numerically found the amplitude of power spectrum on large scale as the function of $h$ and $\delta N$
@@
The amplitude of large-scale curvature perturbation depends on $h$ and $\delta N$ linearly and the amplitude of power spectrum $\Delta_{\zeta}^{\text{(LS)}}\gtrsim 10$, i.e., large-scale amplitude is always greater than the one of single-field inflation.
@@ 4.2 Sharp case
For very small $\delta N$ the Gaussian profile can be approximated to a delta function. For analytically computable, we use the top-hat profile
@@
The wave number $k_{f}\simeq He^{H(t_{1}+t_{2})/2}$ is the scale that the modes cross the horizon during the $h\neq 0$. After the switch-off, the curvature and entropy modes are decoupled. Then the solutions of curvature modes are [55]
@@
where the annihilation and creation operators of two scalar modes satisfying the commutation relations $[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=(2\pi)^{3}\delta_{\alpha\beta}\delta(\boldsymbol{k}-\boldsymbol{k}^{\prime})$ , $[\hat{a}^{\dagger}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}_{\beta}(\boldsymbol{k}^{\prime})]=0$ and $u_{k}$ are the mode functions with Bunch-Davies initial conditons
@@
Then the power spectrum of curvature perurtbation is given
@@
The non-trivial period is the $h_{\text{TH}}\neq 0$ (which is denoted as region-II), where two kinds of modes strongly couple to each other. We also discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
Figure 4: The numerical and analytical results (D) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
@@
We simply use the conformal time as the function of $t$: $\tau(t)=-e^{-Ht}/H$. During $t_{1}$ and $t_{2}$ with $\delta t\equiv t_{2}-t_{1}\ll H^{-1}$, one can discard the Hubble friction of the equation of motions. After $t>t_{1}$, the coupling is switch-on. The duration of $h\neq 0$ is too short $\delta t\ll H^{-1}$ that we can ignore the evolution of the scale factor. Hence we can ignore the Hubble frictions and curvature mass. Then we can use the equations of motion (3.1) and the WKB-type solutions
@@
where the frequencies are given by (3.4). From the equations of motion, we have relations of the coefficients
@@
During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$ hence $p=(k/k_{f})H$ in the above equations. We have eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II and we have eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (4.2) can be solved to (see Appendiex D)
@@
where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$. We show the agreement between numerical and completed analytical results (see appendix D) in Figure 4. The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate during these scales. We found $E_{\zeta}$ and $F_{\zeta}$ are dominated so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
@@
and the rapid oscillation is characterized by $\text{Exp}\left(i2\kappa e^{\pm\delta N/2}\right)$, which is only related to the period of non-zero $h_{\text{tp}}$. The enhancement of power spectrum is provided by low-frequency $\omega_{-}$ and period $\delta N$ together. The scale of peak can be nearly determined by the exponential factor in $\Delta_{\text{env}}$, which is estimated by
@@
These calculations are only valid near scale $k\gtrsim k_{f}$ and for $k\ll k_{f}$ $\Delta_{\zeta}=1$, which miss the matching with the numerical results.
@@
Figure 5: In the top-hat case, the analytical result of large-scale power spectrum as function of $\delta N$ (4), where $h_{0}=24$. The envelope (4.19) is shown as orange dashed curve.
On large scales, the momentum terms in equations of motion can be ignored and the mass terms and all coupling terms become important. The equations of motion can be solved as
@@
We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in the region-II and eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ the coefficients of (4.2) can also be solved [^3] We found that this computation breaks when $x\lesssim 10^{-5}$.. The results are extremely tedious, but we can simple them under some approximations, see appendix D. The results should be real because the perturbations become classical on super-horizon. So we can only consider the real parts
@@
We found that, for fixed $\delta N$, $|G_{\zeta}-H_{\zeta}|$ oscillates around zero when changing $h$. The oscillation is characterized by $\sin(2\sqrt{6}h\delta N)$ and is not significant on amplitude. This is because, on large scales, the dynamics of curvature perturbation is almost a constant and weakly depends on the frequency. For large $h$, the dynamics is determined by an EFT of massless curvature perturbation with sound speed $c_{s}\simeq-1/3$. On the other hand, the amplitudes of $|E_{\zeta}-F_{\zeta}|$ and $|G_{\zeta}-H_{\zeta}|$ exponentially depend on $\delta N$. We can smooth out the oscillation and only estimate the envelope of the power spectrum as
@@
as Figure 5 shows. We can see $\Delta_{\zeta}<1$, in other words, the amplitude of power spectrum on large scales is always smaller than the one of single-field inflation and is not sensitive to $h$ for sharp case. We should mention that this result is also valid for board case with top-hat profile of $h$. For $\delta N\gg 1$, we have $\Delta_{\text{env}}^{\text{(LS)}}\simeq 0.034$, which is not sensitive to $\delta N$ either.

### v3

@@ 2.3 Evolution of the adiabatic mode
where $N_{c}$ is the time of horizon-crossing, $N_{e}$ is the end time of inflation and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing. For small $h$ this is a small correction to the adiabatic mode. The zero-th order entropy perturbation is decoupled with the adiabatic one and massless in the slow-roll limit. Hence the zeroth-order power spectrum of entropy perturbation is simple $\mathcal{P}_{\mathcal{F}}^{(0)}=H^{2}/(4\pi^{2})$. Then the leading-order correction to the power spectrum of curvature perturbation can be estimated to
@@ 3 EFT under massive entropy mode
In this system we have two coupled d.o.f. From the quadratic Lagrangian (2.2) we can write down the linear equations of motion of the system:
@@
Figure 1: Evolution of the entropy perturbation $\mathcal{F}$ for $h=10$. An instability before horizon-crossing and a constant evolution on super-horizon are presented. The scales are arbitrary.
In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator as
@@
The solution of $\mathcal{F}$ is dominated by the $1/m_{s}^{2}$ term, which is given by
@@
We show the consistent of the leading-order contribution of $\mathcal{F}$ in Figure 1 for large enough $h$.
In the slow-roll limit, the entropy mass $m_{s}$ can be calculated by the (2.19). From (B.6) the EFT is only valid under the scale
@@
where $c_{s}$ is the speed of sound of $\zeta$
$\frac{1}{c_{s}^{2}}=1-\frac{4\left(1+2h^{2}\right)}{3+2h^{2}}.$ (3.6)
The operator from the coupling $\dot{\zeta}\mathcal{F}$ would reduce the speed of sound of the fluctuation [45]. We have also used integration by parts to eliminate the $\dot{\zeta}\zeta$ terms in the action. Then we found that the mass term of the curvature fluctuation is cancelled and $\zeta$ is conserved on super-horizon scales. We can find that in the EFT of single field, we have only this constant curvature perturbation on large scales.
Figure 2: The adiabaticity of the super-horizon ($|k\tau_{e}|=10^{-15}$) curvature perturbation. As $h$ increases from very small values, the perturbations become non-adiabatic. However, when $h$ exceeds $\mathcal{O}(0.1)$, the perturbation quickly becomes adiabatic.
[…]
Let’s make some comments on this result:
The evolution of curvature perturbation on large scales is sourced by the non-adiabatic perturbation $\dot{\zeta}=-\delta P_{\text{nad}}H/(\rho+p)$ [47, 48, 49]. Then from the leading-order solution of entropy perturbation, the non-adiabatic mode can be represented as
The adiabatic initial condition ensures that the r.h.s. should be relatively small enough on large scales. We can see that the entropy fluctuation doesn’t decay outside the horizon but is proportional to the curvature fluctuation if $\zeta=\text{constant}$ on the super-horizon. Unlike multi-scalar inflation, where non-decaying non-adiabatic modes will source to curvature modes so that $\zeta$ evolves on the super-horizon. In dilation-gauge inflation, we have constant $\zeta$ for non-decaying $\mathcal{F}$. In other words, For the large $h$ case, $\mathcal{F}$ doesn’t play the role of non-adiabatic mode in dilaton-gauge inflation. We show this non-adiabatic occurrence for mild $h\lesssim\mathcal{O}(0.1)$ in Figure 2.
One of the most important findings from this result is that the cumulative modes (2.26) decay away for large $h$. Because for large $h$, the modes (3.3) on the super-horizon are exactly the constant solutions of the adiabatic mode (2.22). The decaying modes can be also obtained through solving the equation of motion (3), which has been studied in the spatially flat gauge in [20]. The decaying mode is $\zeta\propto(-k\tau)^{(3-\sqrt{9-96h^{2}})/2}$. For small $h$ we can expand this solution and find that the leading correction is proportional to e-folding number $\ln(-k\tau)=N$, which is the solution (2.26).
@@ 4 Primordial black holes
If gauge fields are not important to the inflation, which means that inflation is driven by a single scalar field, it may be ruinous because the higher-order loop corrections to the scalar-field’s mass state indicate that the potential has structure on scales of order of the cut-off of inflation [5]. If the potential has sufficiently steep structure at some points in inflation, i.e., $V_{\phi}>3H\dot{\phi}$ [20], the gauge fields become unstable and can be switched on later. The parameter space for switching on gauge fields is wide making it not difficult for this to occur [50]. We show a simple example in Figure 3. If the dynamics of inflation fall into the parameter space of single-field inflation, the gauge fields will be turned off.
Figure 3: Example of switching gauge fields on during inflation. The model with the potential $V=m^{2}\phi^{2}/2$ and coupling $f=\text{Exp}(c\phi^{2}/M_{\text{pl}}^{2})$, where $m=10^{-6}M_{\text{pl}}$ and $c=3$.
After switching on the gauge fields, the curvature mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound. Therefore the power spectrum is enhanced exponentially. If the vector fields switch on only at late time, the power spectrum is enhanced on small scales, making the formations of PBHs in the radiation era is possible [51, 52]. These PBHs can be candidates for dark matter [53, 54, 55] and supermassive black holes in our universe [56]. The enhancement of the power spectrum due to heavy fields is studied recently in multiple scalar-field models, and is used to discuss PBHs formation [59, 57, 58]. In the previous section, we have realized that it also can be achieved by using vector fields.
We approach to the PBHs without any specific models ($V(\phi)$ and $f(\phi)$). We here only consider the switch-on/off of vector fields at late time $N_{f}$ $only\ once$ and see the features of the power spectrum on small scales for generating PBHs. We choose the horizon-crossing time of the CMB scale $k_{\text{CMB}}=0.05\text{Mpc}^{-1}$ as $N_{\text{CMB}}=0$. Assuming that $h$ has peak at $N_{f}$. The curvature modes with different scales that cross the horizon at different times have different contributions to the power spectrum. We can define the corresponding scale
@@
The modes $k\gg k_{f}$ that cross the horizon far after the switch-off time experienced nothing. Hence the power spectrum is the same as that for single-field inflation $\mathcal{P}_{\zeta}^{(0)}=H^{2}/(8\pi\epsilon M_{\text{pl}}^{2})$ on small scales.
On the other hand, the modes $k\ll k_{f}$ that cross the horizon far before the switch-on time will be impacted by the entropy fluctuation on super-horizon scales. The entropy fluctuation doesn’t decay outside the horizon and thus always transfer to the curvature fluctuation. Hence on these large scales, the power spectrum is shifted and different from that of single-field one.
The interesting scales are those where $k\sim k_{f}$; These modes cross the horizon near the switch-on time $N_{f}$ have large enhancements due to the large instability of the entropy fluctuation. We expect that the power spectrum of the curvature fluctuation has exponential enhancements near the scale $k_{f}$. As shown in Figure 1, the mode takes a few e-folding to grow before freezing outside the horizon. Hence there are two kinds of power spectrum, the broad and the sharp of $h(N)$ [59]. For the sharp case, the curvature modes stop growing before crossing the horizon.
4.1 Broad case
The EFT is valid at time $N_{v}$ satisfying
@@ 4.1 Broad case
Then the majority of contributions to the power spectrum come from modes with $N_{v}\sim N_{f}$. We here consider the Gaussian profile of the function
@@
where the profile is controlled by two parameters $(h_{0},\delta N)$. The "broad" means $\delta N\gtrsim\ln{(h_{0})}$, indicating that we have enough e-folding of non-zero $h(N)$ for exponential growth of the modes. On super-horizon scales, the entropy modes do not decay and still couple to the curvature ones. Hence we should discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
@@
To characterize the peak of the power spectrum for these modes, it’s also useful to introduce a dimensionless parameter $\mathcal{I}$ [46, 60]. The quantization of the curvature fluctuation with imaginary speed of sound differs significantly from that of with real one [46]. The difference lies in the fact that, unlike the Bunch-Davies vacuum, the negative frequency modes are non-zero for initial state of the EFT due to the different commutation conditions of the mode functions. Taking the super-horizon limit, the leading contribution of the the power spectrum is
@@
For modesthat begin to experience transient instability earlier, the amplitude will be larger after crossing the horizon. And we need to use a more UV EFT to describe these modes on deeper sub-horizon scales.
Figure 4: The numerical and analytical results (4.4) of the power spectrum for the Gaussian case, where $h_{0}=8$ and $\delta N=2$.
The scale of the peak $k_{p}$ is the one that has maximal $h$ at the time when the modes are growing at $N_{v}$, i.e., $N_{p_{k}}=N_{v}$. Thus, we have
@@
We present the numerical and analytical results of our calculation in the Figure 4. One point to mention is that (4.4) is not valid for all ranges of $h(N)$ because the computation is under the valid of the EFT, where the mass term of curvature is small compared to that of the heavy entropy (or in other words, for large $h(N)$). Hence, this estimation works better for $|m_{s}^{2}(N)|\gtrsim 8H^{2}$ in our Gaussian-profile model.
@@
As Figure 4 shows, for the l.h.s of the peak, where the switch-on occurs after crossing the horizon, the power spectrum is governed by the long-range contributions. Thus we can observe a shift on large scales due to the coupling with entropy fluctuation on large scales. We have numerically found the amplitude of the power spectrum on large scale as a function of $h$ and $\delta N$
@@
The amplitude of the large-scale curvature perturbation depends linearly on $h$ and $\delta N$, and the amplitude of the power spectrum $\Delta_{\zeta}^{\text{(LS)}}\gtrsim 10$, i.e., the large-scale amplitude is always greater than that of single-field inflation.
@@ 4.2 Sharp case
For very small $\delta N$ the Gaussian profile can be approximated by a delta function. For analytically computation, we use the top-hat profile
@@
The wave number $k_{f}\simeq He^{H(t_{1}+t_{2})/2}$ is the scale at which the modes cross the horizon during the $h\neq 0$. After the switch-off, the curvature and entropy modes decouple. Then the solutions for the curvature modes are [58]
@@
where the annihilation and creation operators of the two scalar modes satisfy the commutation relations $[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=(2\pi)^{3}\delta_{\alpha\beta}\delta(\boldsymbol{k}-\boldsymbol{k}^{\prime})$ , $[\hat{a}^{\dagger}_{\alpha}(\boldsymbol{k}),\hat{a}^{\dagger}_{\beta}(\boldsymbol{k}^{\prime})]=[\hat{a}_{\alpha}(\boldsymbol{k}),\hat{a}_{\beta}(\boldsymbol{k}^{\prime})]=0$ and $u_{k}$ are the mode functions with Bunch-Davies initial conditions
@@
The power spectrum of the curvature perturbation is then given
@@
The non-trivial period is when $h_{\text{TH}}\neq 0$ (which is denoted as region-II), during which the two types of modes strongly couple. We also discuss large-scale modes $k\ll k_{f}$ and small-scale modes $k\sim k_{f}$ separately.
Figure 5: The numerical and analytical results (D) of the power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
@@
We use the conformal time as a function of $t$: $\tau(t)=-e^{-Ht}/H$. During $t_{1}$ and $t_{2}$ with $\delta t\equiv t_{2}-t_{1}\ll H^{-1}$, one can ignore the Hubble friction in the equation of motions. After $t>t_{1}$, the coupling is switch-on. The duration of $h\neq 0$ is too short, $\delta t\ll H^{-1}$, that we can ignore the evolution of the scale factor. Hence we can ignore the Hubble frictions and curvature mass. We then use the equations of motion (B) and the WKB-type solutions:
@@
where the frequencies are given by (B.3). From the equations of motion, we have relations for the coefficients
@@
During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$, hence $p=(k/k_{f})H$ in the above equations. There are eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II, and eight boundary conditions at $t_{1}$, so these coefficients can be completely solved. Then at $t_{2}$, we also have eight boundary conditions and the coefficients of (4.2) can be solved to (see Appendix D)
@@
where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$. We show the agreement between numerical and completed analytical results (see appendix D) in Figure 5. The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate at these scales. We found that $E_{\zeta}$ and $F_{\zeta}$ are dominated, so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
@@
and the rapid oscillation is characterized by $\text{Exp}\left(i2\kappa e^{\pm\delta N/2}\right)$, which is only related to the period of non-zero $h_{\text{tp}}$. The enhancement of the power spectrum is provided by the low-frequency $\omega_{-}$(see Appendix B) and period $\delta N$ together. The scale of the peak can be nearly determined by the exponential factor in $\Delta_{\text{env}}$, which is estimated by
@@
These calculations are only valid near the scale $k\gtrsim k_{f}$ and for $k\ll k_{f}$ $\Delta_{\zeta}=1$, which miss the matching with the numerical results.
@@
Figure 6: In the top-hat case, the analytical result of the large-scale power spectrum as a function of $\delta N$ (4), where $h_{0}=24$. The envelope (4.19) is shown as the orange dashed curve.
On large scales, the momentum terms in the equations of motion can be ignored, and the mass terms and all coupling terms become important. The equations of motion can be solved as
@@
We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in region-II and eight boundary conditions at $t_{1}$, so all of these coefficients can be completely solved. Then at $t_{2}$ the coefficients of (4.2) can also be solved [^4] We found that this computation breaks when $x\lesssim 10^{-5}$.. The results are extremely tedious, but we can simplify them under some approximations, see appendix D. The results should be real because the perturbations become classical on the super-horizon. So we can only consider the real parts
@@
We found that, for fixed $\delta N$, $|G_{\zeta}-H_{\zeta}|$ oscillates around zero when changing $h$. The oscillation is characterized by $\sin(2\sqrt{6}h\delta N)$ and is not significant in amplitude. This is because, on large scales, the dynamics of the curvature perturbation is almost constant and weakly depends on the frequency. For large $h$, the dynamics is determined by an EFT of massless curvature perturbation with sound speed $c_{s}\simeq-1/3$. On the other hand, the amplitudes of $|E_{\zeta}-F_{\zeta}|$ and $|G_{\zeta}-H_{\zeta}|$ exponentially depend on $\delta N$. We can smooth out the oscillation and only estimate the envelope of the power spectrum as
@@
as Figure 6 shows. We can see $\Delta_{\zeta}<1$, in other words, the amplitude of the power spectrum on large scales is always smaller than that of single-field inflation and is not sensitive to $h$ for sharp case. We should mention that this result is also valid for broad case with a top-hat profile of $h$. For $\delta N\gg 1$, we have $\Delta_{\text{env}}^{\text{(LS)}}\simeq 0.034$, which is not sensitive to $\delta N$ either.